OpenStudy (anonymous):

When you solve an equation, how do you know when to add and when to subtract? You may include an example if you need to.

OpenStudy (anonymous):

y-10=2y+5 I would add 10 to both sides

OpenStudy (anonymous):

\begin{array}l\color{red}{\normalsize\text{U}}\color{orange}{\normalsize\text{ }}\color{#9c9a2e}{\normalsize\text{W}}\color{green}{\normalsize\text{i}}\color{blue}{\normalsize\text{n}}\color{purple}{\normalsize\text{ }}\color{purple}{\normalsize\text{M}}\color{red}{\normalsize\text{o}}\color{orange}{\normalsize\text{N}}\color{#9c9a2e}{\normalsize\text{e}}\color{green}{\normalsize\text{Y}}\color{blue}{\normalsize\text{!}}\color{purple}{\normalsize\text{!}}\color{purple}{\normalsize\text{!}}\color{red}{\normalsize\text{ }}\color{orange}{\normalsize\text{}}\end{array}

OpenStudy (anonymous):

@ Daveboyback: Are you still here? I can explain if you promise to focus here for a couple of minutes.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Awesome. Whenever you solve an equation, you usually want it to be in the form of y = something. I'll check on each step to make sure that everything makes sense. So far so good, yes?

OpenStudy (anonymous):

yes'

OpenStudy (anonymous):

Awesome. Now give me a random example that has X's and a Y in them and we'll use it as an example to solve.

OpenStudy (anonymous):

Also, make sure the y is not by itself yet, so that we can actually do work to it.

OpenStudy (anonymous):

yeah im sory but i kind of finished this but umm what i realy need is this When you solve an equation, how do you know when to add and when to subtract? You may include an example if you need to.

OpenStudy (anonymous):

That's what we're doing. :P I'll go ahead and make an example then. But first: have you done equations with both x and y in them? Or did they only teach you equations with just X in them?

OpenStudy (anonymous):

ohh we are i thought i was in me other quesion!!

OpenStudy (anonymous):

Ok, cool. So does something like 2y+3x + 5 = 10 look familiar?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Cool! So how do we want our equation to look again? Recall what I said waaayyyy earlier about "we always want our equations to look like __ = ___"

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Just scroll up and tell me the first thing we agreed to. I said we were going to go step by step to make sure you understand everything and so far you agreed to one thing. So see what we agreed to.

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Alright, so what do we want every equation to look like?